# VTAO-BiManip: Masked Visual-Tactile-Action Pre-training with Object Understanding for Bimanual Dexterous Manipulation

### v1

@@ IV-D Real World Experiment
We create a real-world experimental platform with dual Leap Hands [2] and a Unitree Z1 arm  [^3] https://www.unitree.com/z1, as shown in Fig. 6. The simulation task for bottle cap unscrewing is recreated in Isaac Gym [49] using dual Leap Hands. The final validation confirms that our method supports effective physical experimentation.
@@ V Conclusion
This paper investigates the effectiveness of a pre-trained representation model integrating multi-modalities, action prediction, and object understanding for dexterous bimanual manipulation. We design a novel pre-training framework, VTAO-BiManip, which is pre-trained using human VTAO demonstrations. We integrate a learning curriculum into RL to address the challenge of multi-skill learning within a bimanual dexterous skill. The representation model is incorporated into the curriculum RL framework for bimanual manipulation. Experimental results show that our approach significantly outperforms baseline methods. This study represents a pioneering effort in multimodal bimanual manipulation. Future work may focus on: 1) developing large-scale multimodal bimanual datasets across diverse scenes and tasks; 2) exploring additional bimanual manipulation tasks; 3) leveraging unpaired existing 3D object labeling data.

### v2

@@ IV-D Real World Experiment
We create a real-world experimental platform with dual Leap Hands [shaw2023leaphand] equipped with tactile sensors and a Unitree Z1 arm  [^3] https://www.unitree.com/z1, as shown in Fig. 6. Visual perception is obtained through a ego-centric RGB camera, while tactile signals is captured via tactile sensors mounted on the dual Leap Hands and processed through binarization. The simulation task for bottle cap unscrewing is recreated in Isaac Gym [makoviychuk2021isaac] using dual Leap Hands. The final validation confirms that our method supports effective physical experimentation. Please refer to our submitted video for detailed demonstrations of the system’s real-world performance.
@@ V Conclusion
This paper investigates the effectiveness of a pre-trained representation model integrating multi-modal input, action prediction, and object understanding for dexterous bimanual manipulation. We design a novel pre-training framework, VTAO-BiManip, which is pre-trained using human VTAO demonstrations. We integrate a learning curriculum into RL to address the challenge of multi-skill learning within a bimanual dexterous skill. The representation model is incorporated into the curriculum RL framework for bimanual manipulation. Experimental results show that our approach significantly outperforms baseline methods. This study represents a pioneering effort in multimodal bimanual manipulation. Future work may focus on: 1) Scaling multimodal bimanual datasets across diverse scenes and tasks; 2) Extending to broader manipulation scenarios; 3) Exploring geometric object priors from vision models. Our method utilizes motion-captured object 3D bounding boxes (pose+size) for the object understanding. However, it may be replaced by poses estimated by large models for object pose estimaiton in the future.